OpenStudy (anonymous):

Find all critical values for the function f(x)=x^1/13(x−1)^2

OpenStudy (anonymous):

not clear what function is. looks like \[f(x)=\frac{f^1}{13(x-1)^2}\] but that seems unlikely. it is \[f(x)=x^{\frac{1}{13}}\times (x-1)^2\]?